Suppose z is a function of x and y and tan(√x + y) = e^z^2. Determine ∂z/∂x and ∂z/∂y.

1
Expert's answer
2022-05-30T02:28:51-0400
"\tan(\sqrt{x}+y)=e^{z^2}"

Differentiate both sides with respect to "x"


"\dfrac{1}{\cos ^2(\sqrt{x}+y)}(\dfrac{1}{2\sqrt{x}})=2ze^{z^2}\dfrac{\partial z}{\partial x}"

"\dfrac{\partial z}{\partial x}=\dfrac{1}{4ze^{z^2}\sqrt{x}\cos ^2(\sqrt{x}+y)}"



"\tan(\sqrt{x}+y)=e^{z^2}"

Differentiate both sides with respect to "y"

"\dfrac{1}{\cos ^2(\sqrt{x}+y)}(1)=2ze^{z^2}\dfrac{\partial z}{\partial y}"

"\dfrac{\partial z}{\partial y}=\dfrac{1}{2ze^{z^2}\cos ^2(\sqrt{x}+y)}"